A man is 24 years older than his son. In two years, his age will be twice the age of his son. what is the present age of his son?
Answer by mananth(16949)

A man is 24 years older than his son. In two years, his age will be twice the age of his son. what is the present age of his son?
...
son's age = xyears
father's age = x=24 years
..
after 2 years
son = x+2
father = x+24+2= x+26
..
x+26=2(x+2)
x+26=2x+4
2x-x=26-4
x=22 years son's age
father's age = 24+22=46 years
..
m.ananth@hotmail.ca